OpenStudy (anonymous):

If the measures of the angles of a triangle are in the ratio of 1:2:3, then the lengths of the sides are in the ratio of 1:2:3. True False

OpenStudy (anonymous):

will definitely award medals

OpenStudy (phi):

do you know the Law of Sines? it says for sides a and b and opposite angles A and B: \[ \frac{b}{\sin(B)}= \frac{a}{\sin(A)} \] or, if we rearrange \[ \frac{\sin(A)}{\sin(B)}= \frac{a}{b} \] if the ratio of sides a/b is ½ then we have \[ \frac{\sin(A)}{\sin(B)}= \frac{1}{2} \] notice the ratio of angles A/B is not ½ it is the ratio of the sin(A) to sin(B) that is 1/2

OpenStudy (anonymous):

so is it true or false

OpenStudy (anonymous):

no i don't know th law od sines

OpenStudy (phi):

did you read what I posted ? It tells you the answer if you think about it.

OpenStudy (phi):

one reason to study math is to force people to think.

OpenStudy (anonymous):

false right?

OpenStudy (phi):

notice the ratio of angles A/B is not ½ the ratio of the angles is not the same as the ratio of the sides

OpenStudy (anonymous):

so the answer is false

OpenStudy (phi):

yes, it is false